# Natural solution of SUSY $\mu$ problem from modulus stabilization in modular flavor model

### v3

@@ Abstract
We propose a solution to the SUSY $\mu$-problem within the framework of modular flavor symmetry. The explicit $\mu$-term is prohibited by modular symmetry, and an effective $\mu$-term is regenerated following the stabilization of the modulus field. We examine the stabilization mechanism of a single modulus field with the presence of SUSY breaking contributions described by the non-linear SUSY realization scheme involving a nilpotent Goldstino $\textbf{X}_{nl}$ superfield. The inclusion of $\textbf{X}_{nl}$ contributions can slightly alter the original stabilized modulus VEV to a value near some fixed point. A natural small $\mu_{eff}$, significantly smaller than the SUSY scale, can result from either the expansion of typical modular forms using a small deviation parameter near the fixed point $\omega$, or from the combined effects of suppression by powers of $q^{1/24}$ [or $(2\Im\tau)^{-1}$] along with the asymptotic suppression behavior of typical modular forms away from the fixed point $i\infty$, taking the form of appropriate power of the tiny deviation parameter.
@@ 1 Introduction
Modular symmetry is a geometrical symmetry of torus $T^{2}$ and it corresponds to a change of the corresponding cycle basis. Modular symmetry can naturally arise in string theory, for example, in the (exact) moduli spaces of Calabi-Yau threefolds like the quintic, and more generically in large classes of elliptic Calabi-Yau threefolds. Recently, modular flavor symmetry was proposed in Feruglio:2017spp to explain the flavor puzzles of the standard model (plus the massive neutrinos), which no longer necessitates the introduction of the flavon fields that may cause difficulties in flavor model buildings. For reviews, see Feruglio:2019ybq ; Kobayashi:2023zzc ; Ding:2023htn ; Ding:2024ozt . In the modular flavor symmetry framework, the superpotential is tightly constrained by the modular symmetry, with the relevant Yukawa couplings being modular forms of a certain level $N$, which are functions of a modulus field. The matter superfields and the modular forms in the superpotential transform in some representations of certain finite modular group $\Gamma_{N}\equiv\bar{\Gamma}/\bar{\Gamma}(N),N\leq 5$ (or its double cover $\Gamma_{N}^{\prime}$). In this framework, the masses and mixings of quarks and leptons can be predicted with very few input parameters. Modular flavor models based on $\Gamma_{2}\simeq S_{3}$ Kobayashi:2018vbk ; Okada:2019xqk ; Du:2020ylx , $\Gamma_{3}\simeq A_{4}$ Kobayashi:2018vbk ; Petcov:2018snn ; Criado:2018thu ; Kobayashi:2018scp ; Okada:2018yrn ; Novichkov:2018yse ; Okada:2020ukr ; Petcov:2022fjf ; CentellesChulia:2023osj ; Kumar:2023moh ; Nomura:2024ctl ; Pathak:2024sei , $\Gamma_{4}\simeq S_{4}$ Penedo:2018nmg ; Petcov:2018snn ; Novichkov:2018ovf ; Kobayashi:2019xvz ; Wang:2020dbp ; Qu:2021jdy , $\Gamma_{5}\simeq A_{5}$ Petcov:2018snn ; Novichkov:2018nkm ; Ding:2019xna ; Yao:2020zml ; deMedeirosVarzielas:2022ihu ; Abbas:2024bbv and non-holomorphic modular flavor models Qu:2024rns ; Nomura:2024atp ; Ding:2024inn ; Li:2024svh had been used in various studies to explain the low energy flavor structures of the Standard Model. Modular symmetry can also be used to solve the strong CP problem Feruglio:2023uof ; Feruglio:2024ytl . Moreover, modular flavor symmetries can be combined with the GUT framework to further enhance the predictive power of the models, such as combining $\Gamma_{2},\Gamma_{3}$ with $SU(5)$ deAnda:2018ecu ; Kobayashi:2019rzp ; Du:2020ylx ; Chen:2021zty ; Zhao:2021jxg ; King:2021fhl ; Ding:2021zbg , SO(10) Ding:2021eva ; Ding:2022bzs or flipped SU(5) Charalampous:2021gmf ; Du:2022lij ; King:2024gsd GUT models.
@@
The vacuum expectation value (VEV) of the modulus field can be the only source of flavour symmetry breaking. It can also be the only source of CP symmetry breaking Novichkov:2019sqv ; Feruglio:2021dte ; Feruglio:2024ytl . The VEV of the modulus field $\tau$ generally break the modular symmetry completely, except at three inequivalent fixed points $\tau=i\infty,i,e^{i2\pi/3}$ in the fundamental domain. The modular forms (thus the Yukawa couplings) and the fermion mass matrices are fixed once the modular flavour symmetry is fully or partially broken by the VEV of $\tau$. From a top-down perspective, the value of modulus VEV can be determined dynamically with a modular-invariant scalar potential in 4d $N=1$ supergravity. This is also crucial for giving nonzero masses to the modulus field, rendering it phenomenological acceptable.
@@
Experiences from heterotic string theory on moduli stabilization target:space ; Leedom:2022zdm ; Gonzalo:2018guu indicate that the scalar potential favors minima at the boundary of the fundamental domain or at points inside the fundamental domain that are close to the fixed point $\omega=e^{i2\pi/3}$ with negative vacuum energy. In the presence of additional ”matter”-like fields, it is possible to obtain a dS vacua with additional uplifting contributions, such as the introduction of an anti-D3 brane that provides a positive uplifting potential, which would induce a SUSY breaking vacuum shift to a tiny positive vacuum energy. The presence of the anti-brane-like sources can be described in terms of a nilpotent Goldstino superfield KS in a supersymmetric effective field theory. In the framework of non-linearly realized SUSY, the couplings of chiral, gauge and gravity superfields to the constrained nilpotent Goldstino superfield $X$ can be used to describe spontaneous SUSY breaking in supergravity Burgess:2021juk ; Aparicio:2015psl ; NL:SUGRA1 ; NL:SUGRA2 ; NL:SUGRA3 ; NL:SUGRA4 ; NL:SUGRA5 ; NL:SUGRA6 and the presence of the anti-brane type sources with a supersymmetric action. Contributions from nilpotent chiral superfield can uplift the scalar potential responsible for the stabilization of the modulus field.
@@
This paper is organized as follows. In Sec 2, we briefly review the $\Gamma_{N}$ modular flavor symmetry. In Sec 3, we discuss the solution of SUSY $\mu$-problem using modular flavor symmetry. In Sec 4, we discuss the stabilization of singlet modulus field in the presence of SUSY breaking contributions described by the non-linear realized SUSY scheme with the nilpotent Goldstino ${\textbf{X}}_{nl}$ superfield. In Sec 5, we discuss the regeneration of a naturally small $\mu_{eff}$ after modulus stabilization, when the modulus field is stabilized near some fixed points, such as $\omega\equiv e^{i2\pi/3}$ and $i\infty$ Sec 6 contains our conclusions.
@@ 2 Modular $\Gamma_{N}$ symmetry
$S:\tau\mapsto-\frac{1}{\tau},~{}~{}~{}~{}\quad T:\tau\mapsto\tau+1\,,$ (2)
@@
$S=\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right),~{}~{}~{}\quad T=\left(\begin{array}[]{cc}1&1\\ 0&1\end{array}\right)\,.$ (4)
@@
with $\Gamma(N)$ the principal congruence subgroup of level $N$
@@ 3 Natural $\mu$-term in modular flavor model
$+$ $\int d^{2}\theta\left[\alpha\Lambda_{SUSY}\frac{Y_{N}^{k_{r}}(\tau)}{\eta(\tau)^{2k_{3}}}H_{u}H_{d}+\cdots\right],$
@@
$\tilde{\mu}=\alpha\Lambda_{SUSY}\frac{Y_{N}^{(k_{r})}(\tau)}{\eta(\tau)^{2k_{3}}}~{},$ (13)
@@
where $\epsilon_{\eta}(\gamma)$ depends on $\gamma$ (but not on $\tau$) and can be expressed in terms of the Rademacher $\Phi(\gamma)$ function
$\epsilon_{\eta}(\gamma)=\exp\left[\frac{2\pi i}{24}\Phi(\gamma)-i\frac{\pi}{4}\right],$ (16)
that satisfies $\epsilon_{\eta}(\gamma)^{24}=1$ DHoker:2022dxx .
@@
$\mu_{eff}=\alpha\Lambda_{SUSY}\frac{Y_{N}^{k_{r}}(\tau_{0})}{\eta(\tau_{0})^{2k_{3}}}(2\Im\tau_{0})^{(k_{H_{u}}+k_{H_{d}})/2}=\alpha\Lambda_{SUSY}\frac{Y_{N}^{k_{r}}(\tau_{0})}{\eta(\tau_{0})^{2k_{3}}}(2\Im\tau_{0})^{(k_{r}-k_{3}+3)/2}.$ (19)
We need to explain why $\mu_{eff}$ is much smaller than $\Lambda_{SUSY}$ and of order the electroweak scale or soft SUSY breaking scale $m_{soft}$. We will show that, the smallness of $\mu_{eff}$ is the consequence of large suppression factors, either originate from some power of deviation parameter that characterizes the deviation of modulus VEV from typical fixed point, or from non-trivial power (determined by the choice of modular weights for $H_{u}$ and $H_{d}$) of a small factor related to the modulus VEV. All these suppression factors are determined by the modulus stabilization mechanism, with the stabilized modulus VEVs being affected by SUSY breaking effects.
@@ 4 Stabilization of modulus field with nilpotent Goldstino ${\textbf{X}}_{nl}$ superfield
To stabilize the modulus field, we can utilize effective supergravity scalar potentials derived from a weight $-3$ superpotential to determine the minimum. The modulus VEV can always be restricted to the fundamental domain
@@
$\equiv$ $\left(\frac{G_{6}(\tau)}{\eta(\tau)^{12}}\right)^{m}\left(\frac{G_{4}(\tau)}{\eta(\tau)^{12}}\right)^{n}{\cal P}\left(j(\tau)\right)~{},$
where ${\cal P}\left(j(\tau)\right)$ denotes a polynomial with respect to $j(\tau)$. The Klein $j$-function is defined as a modular form of zero weight
$j(\tau)$$\equiv$ $\frac{3^{6}5^{3}}{\pi^{12}}\frac{G_{4}(\tau)^{3}}{\eta(\tau)^{24}},$ (25)
@@
with $j(i\infty)$ diverges, $j(\omega)=0$ and $j(i)=1728$.
@@
where all the fields are the functions of $y^{\mu}\equiv x^{\mu}+i\theta\sigma^{\mu}\bar{\theta}$. The lowest component of $\textbf{X}_{nl}$ is eliminated by the nilpotent constraints $\textbf{X}_{nl}^{2}=0$ and is no longer an independent complex scalar field. This leaves only the Goldstino field $\psi_{X}$ and an independent auxiliary field $F_{X}$ in $\textbf{X}_{nl}$ superfield. The Goldstino field $\psi_{X}$, which is the fermionic component of $\textbf{X}_{nl}$, eventually mixes with the gravitino to give it a mass through the super-Higgs mechanism. The lowest component of $\textbf{X}_{nl}$, which is denoted as ${X}_{nl}$ for the Goldstino bilinear ${\psi_{X}\psi_{X}}$, obviously satisfies ${X}_{nl}^{2}=0$ due to its anti-commuting Grassman variable nature. The chiral superfield $\textbf{X}_{nl}$ can be seen as the long-distance expression of the microscopic superfield X that satisfies $\bar{D}^{\dot{\alpha}}{\cal J}_{\alpha\dot{\alpha}}=D_{\alpha}\textbf{X}$ for Ferrara-Zumino multiplet ${\cal J}_{\alpha\dot{\alpha}}$ KS , which can be used even when SUSY is spontaneously broken and can be followed consistently along the renormalization group flow down to the deep IR. As noted in Aparicio:2015psl , for the anti-D3 brane in the KKLT scenario, the representation in terms of $\textbf{X}_{nl}$ is very convenient and enough to capture all the anti-D3 brane degrees of freedoms since it allows to treat the effect of anti-D3 brane in terms of standard supergravity couplings of matter and moduli superfields to the nilpotent Goldstino superfield. From the perspective of low energy effective field theory, regardless of the UV origin of SUSY breaking, such a non-linear realization scheme is also convenient for discussions on modulus stabilization and vacuum uplifting.
@@
[…]
@@
where without loss of generality we take $f_{0}$ and $c_{1}$ to be real. Integrating out the auxillary field $F_{X}$ gives $F_{X}=-(W_{\textbf{X}_{nl}})^{*}=-f_{0}$. To accommodate the modular flavor symmetry, the Kahler potential and superpotential needs to be adapted to
@@
$W$ $\supseteq$ $\left(f_{0}\textbf{X}_{nl}\left[\eta(\tau)\right]^{2k_{X}}+c_{1}^{3}\right)\frac{H(\tau)}{\eta^{6}(\tau)}.$ (30)
The supergravity scalar potential can be derived from
@@
$K^{{\alpha}\bar{\beta}}$ $=$ $\left(\begin{array}[]{cc}\frac{(2\Im\tau)^{2}}{3}\left[1-\frac{k_{X}}{3}\frac{\textbf{X}^{\dagger}_{nl}\textbf{X}_{nl}}{(2\Im\tau)^{k_{X}}}\right]&\frac{i(2\Im\tau)k_{X}\textbf{X}_{nl}}{3}\\ -\frac{i(2\Im\tau)k_{X}\textbf{X}^{\dagger}_{nl}}{3}&(2\Im\tau)^{k_{X}}\left[1+\frac{k_{X}^{2}(\textbf{X}^{\dagger}_{nl}\textbf{X}_{nl})}{3(2\Im\tau)^{k_{X}}}\right]\end{array}\right)~{}.$ (35)
Here the indices $\alpha,\beta\sim(\tau,\textbf{X}_{nl})$ and the nilpotency of $\textbf{X}_{nl}$ is used in the calculation.
The scalar potential for the modulus field $\tau$ and $X_{nl}$ can be calculated as
@@
-- $\left.3\left|\tilde{c}_{W}\right|^{2}\left|{H(\tau)}\right|^{2}\right\}~{},$
@@
$\tilde{W}_{T}$$\equiv$ $k_{X}f_{0}X_{nl}\left[\eta(\tau)\right]^{2k_{X}}\left[\frac{i}{2\pi}\hat{G}_{2}(\tau,\tau^{*})+\frac{i}{2\Im\tau}+\frac{ic_{1}^{3}X_{nl}^{*}}{f_{0}(2\Im\tau)^{k_{X}+1}\eta(\tau)^{2k_{X}}}\right]~{},$
$\tilde{c}_{M}$ $=$ $f_{0}X_{nl}\left[\eta(\tau)\right]^{2k_{X}}+c_{1}^{3}~{},$
$\tilde{W}_{X}$ $=$ $f_{0}\eta(\tau)^{2k_{X}}\left(1+\frac{|X_{nl}|^{2}}{2(\Im\tau)^{k_{X}}}\right)+\frac{{X}^{*}_{nl}c_{1}^{3}}{(2\Im\tau)^{k_{X}}}.$ (37)
The constraint $X_{nl}^{2}=0$ for the lowest component of $\textbf{X}_{nl}$ is used to eliminate some of the high order $X_{nl}$ terms.
We also use the following expression for the derivative of Dedekind $\eta$ function
$\frac{\eta^{\prime}(\tau)}{\eta(\tau)}=\frac{i}{4\pi}G_{2}~{},$ (38)

### v4

@@ Abstract
We propose a solution to the SUSY $\mu$-problem within the framework of modular flavor symmetry. The explicit $\mu$-term is prohibited by modular symmetry, and an effective $\mu$-term is regenerated following the stabilization of the modulus field. We examine the stabilization mechanism of a single modulus field with the presence of SUSY breaking contributions described by the non-linear SUSY realization scheme involving a nilpotent Goldstino $\textbf{X}_{nl}$ superfield. A natural small $\mu_{eff}$, significantly smaller than the SUSY scale, can result from either the expansion of typical modular forms using a small deviation parameter near the fixed point $\omega$, or from the combined effects of suppression by powers of $q^{1/24}$ [or $(2\Im\tau)^{-1}$] along with the asymptotic suppression behavior of typical modular forms away from the fixed point $i\infty$, taking the form of appropriate power of the tiny deviation parameter. A natural small $\mu_{eff}$ can also be achieved by a weighton-like mechanism for $H_{u}H_{d}$ bilinear.
@@ 1 Introduction
Modular symmetry is a geometrical symmetry of torus $T^{2}$ and it corresponds to a change of the corresponding cycle basis. Modular symmetry can naturally arise in string theory, for example, in the (exact) moduli spaces of Calabi-Yau threefolds like the quintic, and more generically in large classes of elliptic Calabi-Yau threefolds CY:modular1 ; CY:modular2 . Recently, modular flavor symmetry was proposed in Feruglio:2017spp to explain the flavor puzzles of the Standard Model (plus the massive neutrinos), which no longer necessitates the introduction of the flavon fields that may cause difficulties in flavor model buildings. For reviews, see Feruglio:2019ybq ; Kobayashi:2023zzc ; Ding:2023htn ; Ding:2024ozt . In the modular flavor symmetry framework, the superpotential is tightly constrained by the modular symmetry, with the relevant Yukawa couplings being modular forms of a certain level $N$, which are functions of a modulus field. The matter superfields and the modular forms in the superpotential transform in some representations of certain finite modular group $\Gamma_{N}\equiv\bar{\Gamma}/\bar{\Gamma}(N),N\leq 5$ (or its double cover $\Gamma_{N}^{\prime}$). In this framework, the masses and mixings of quarks and leptons can be predicted with very few input parameters. Modular flavor models based on $\Gamma_{2}\simeq S_{3}$ Kobayashi:2018vbk ; Okada:2019xqk ; Du:2020ylx , $\Gamma_{3}\simeq A_{4}$ Kobayashi:2018vbk ; Petcov:2018snn ; Criado:2018thu ; Kobayashi:2018scp ; Okada:2018yrn ; Novichkov:2018yse ; Okada:2020ukr ; Petcov:2022fjf ; CentellesChulia:2023osj ; Kumar:2023moh ; Nomura:2024ctl ; Pathak:2024sei , $\Gamma_{4}\simeq S_{4}$ Penedo:2018nmg ; Petcov:2018snn ; Novichkov:2018ovf ; Kobayashi:2019xvz ; Wang:2020dbp ; Qu:2021jdy , $\Gamma_{5}\simeq A_{5}$ Petcov:2018snn ; Novichkov:2018nkm ; Ding:2019xna ; Yao:2020zml ; deMedeirosVarzielas:2022ihu ; Abbas:2024bbv and non-holomorphic modular flavor models Qu:2024rns ; Nomura:2024atp ; Ding:2024inn ; Li:2024svh had been studied in various works to explain the low energy flavor structures of the Standard Model. Modular symmetry can also be used to solve the strong CP problem Feruglio:2023uof ; Feruglio:2024ytl . Moreover, modular flavor symmetries can be combined with the GUT framework to further enhance the predictive power of the models, such as combining $\Gamma_{2},\Gamma_{3}$ with $SU(5)$ deAnda:2018ecu ; Kobayashi:2019rzp ; Du:2020ylx ; Chen:2021zty ; Zhao:2021jxg ; King:2021fhl ; Ding:2021zbg , SO(10) Ding:2021eva ; Ding:2022bzs or flipped SU(5) Charalampous:2021gmf ; Du:2022lij ; King:2024gsd GUT models.
@@
The vacuum expectation value (VEV) of the modulus field can be the only source of flavour symmetry breaking. It can also be the only source of CP symmetry breaking Baur:2019kwi ; Novichkov:2019sqv ; Feruglio:2021dte ; Ishiguro:2020nuf ; Feruglio:2024ytl . The VEV of the modulus field $\tau$ generally break the modular symmetry completely, except at three inequivalent fixed points $\tau=i\infty,i,e^{i2\pi/3}$ in the fundamental domain. The modular forms (thus the Yukawa couplings) and the fermion mass matrices are fixed once the modular flavour symmetry is fully or partially broken by the VEV of $\tau$. From a top-down perspective, the value of modulus VEV can be determined dynamically with a modular-invariant scalar potential in 4d $N=1$ supergravity. This is also crucial for giving nonzero masses to the modulus field, rendering it phenomenological acceptable.
@@
Experiences from heterotic string theory on moduli stabilization target:space ; Leedom:2022zdm ; Gonzalo:2018guu indicate that the scalar potential favors minima at the boundary of the fundamental domain or at points inside the fundamental domain that are close to the fixed point $\omega=e^{i2\pi/3}$ with negative vacuum energy. In the presence of additional ”matter”-like fields, it is possible to obtain a dS vacua with additional uplifting contributions, such as the introduction of an anti-D3 brane that provides a positive uplifting potential, which would induce a SUSY breaking vacuum shift to a tiny positive vacuum energy. The presence of the anti-brane-like sources can be described in terms of a nilpotent Goldstino superfield KS in a supersymmetric effective field theory. In the framework of non-linearly realized SUSY, the couplings of chiral, gauge and gravity superfields to the constrained nilpotent Goldstino superfield $\textbf{X}_{nl}$ can be used to describe spontaneous SUSY breaking in supergravity Burgess:2021juk ; Aparicio:2015psl ; NL:SUGRA1 ; NL:SUGRA2 ; NL:SUGRA3 ; NL:SUGRA4 ; NL:SUGRA5 ; NL:SUGRA6 and the presence of the anti-brane type sources with a supersymmetric action. Contributions from nilpotent chiral superfield can uplift the scalar potential responsible for the stabilization of the modulus field.
@@
This paper is organized as follows. In Sec 2, we briefly review the $\Gamma_{N}$ modular flavor symmetry. In Sec 3, we discuss the solution of SUSY $\mu$-problem using modular flavor symmetry. In Sec 4, we discuss the stabilization of singlet modulus field in the presence of SUSY breaking contributions described by the non-linear realized SUSY scheme with the nilpotent Goldstino ${\textbf{X}}_{nl}$ superfield. In Sec 5, we discuss the regeneration of a naturally small $\mu_{eff}$ after modulus stabilization, when the modulus field is stabilized near some fixed points, such as $\omega\equiv e^{i2\pi/3}$ and $i\infty$ . Sec 6 contains our conclusions.
@@ 2 Modular $\Gamma_{N}$ symmetry
$S:\tau\mapsto-\frac{1}{\tau},~~~~\quad T:\tau\mapsto\tau+1\,.$ (2)
@@
$S=\left(\begin{array}[]{cc}0&1\\ -1&0\end{array}\right),~~~\quad T=\left(\begin{array}[]{cc}1&1\\ 0&1\end{array}\right)~.$ (4)
@@
with $\Gamma(N)$ the principal congruence subgroup of level NN
@@ 3 Natural $\mu$-term in modular flavor model
$+$ $\int d^{2}\theta\left[\alpha\Lambda_{SUSY}\frac{Y_{N}^{(k_{r})}(\tau)}{\left[\eta(\tau)\right]^{2k_{3}}}H_{u}H_{d}+\cdots\right],$
@@
$\tilde{\mu}=\alpha\Lambda_{SUSY}\frac{Y_{N}^{(k_{r})}(\tau)}{\left[\eta(\tau)\right]^{2k_{3}}}~,$ (13)
@@
where $\epsilon_{\eta}(\gamma)$ depends on $\gamma$ (but not on $\tau$) and can be expressed as
$\epsilon_{\eta}(\gamma)$ $=$ $\exp\left[\frac{\pi i}{12}\omega(a,b,c,d)-i\frac{\pi}{4}\right],$~ (16)
that satisfies $\epsilon_{\eta}(\gamma)^{24}=1$ DHoker:2022dxx . Here
$\omega(a,b,c,d)=\left(\frac{a+d}{c}+12s(-d,c)\right),$ (17)
is an integer and s(d,c)s(-d,c) is a Dedekind sum.
@@
$\mu_{eff}=\alpha\Lambda_{SUSY}\frac{Y_{N}^{(k_{r})}(\tau_{0})}{\left[\eta(\tau_{0})\right]^{2k_{3}}}(2\Im\tau_{0})^{(k_{H_{u}}+k_{H_{d}})/2}=\alpha\Lambda_{SUSY}\frac{Y_{N}^{(k_{r})}(\tau_{0})}{\left[\eta(\tau_{0})\right]^{2k_{3}}}(2\Im\tau_{0})^{(k_{r}-k_{3}+3)/2}.$ (20)
We need to explain why $\mu_{eff}$ is much smaller than $\Lambda_{SUSY}$ and of order the electroweak scale or soft SUSY breaking scale $m_{soft}$. We will show that, the smallness of $\mu_{eff}$ is the consequence of large suppression factors, either originate from some power of deviation parameter that characterizes the deviation of modulus VEV from typical fixed point, or from non-trivial power (determined by the choice of modular weights for $H_{u}$ and $H_{d}$) of a small factor related to the modulus VEV. All these suppression factors are determined by the modulus stabilization mechanism, with the stabilized modulus VEVs being affected by SUSY breaking effects. The smallness of $\mu_{eff}$ can also be achieved by a weighton-like mechanism King:2020qaj for $H_{u}H_{d}$ bilinear.
@@ 4 Stabilization of modulus field with nilpotent Goldstino ${\textbf{X}}_{nl}$ superfield
To stabilize the modulus field, we can utilize effective supergravity scalar potentials derived from a weight 3-3 superpotential to determine the minimum. The modulus VEV can always be restricted to the fundamental domain
@@
$\equiv$ $\left(\frac{G_{6}(\tau)}{\left[\eta(\tau)\right]^{12}}\right)^{m}\left(\frac{G_{4}(\tau)}{\left[\eta(\tau)\right]^{8}}\right)^{n}{\cal P}\left(j(\tau)\right)~,$
where ${\cal P}\left(j(\tau)\right)$ denotes a polynomial with respect to $j(\tau)$ and $m,n$ are non-negative integers target:space . The definitions of $G_{4}(\tau)$ and $G_{6}(\tau)$ are given in the Appendix B. The Klein $j$-function is defined as a modular form of zero weight
$j(\tau)$ $\equiv$ $\frac{3^{6}5^{3}}{\pi^{12}}\frac{G_{4}(\tau)^{3}}{\left[\eta(\tau)\right]^{24}},$ (26)
@@
with $j(i\infty)$ diverges, $j(\omega)=0$ and $j(i)=1728$. Note that $j(\tau)$ has a triple zero at $\tau=\omega$, and $j(\tau)-1728$ has a double zero at $\tau=i$. The derivative $j^{\prime}(\tau)$ vanishes only at these values: $j^{\prime}(\omega)=j^{\prime}(i)=0$ Klein:jfunction .
@@
where all the fields are the functions of $y^{\mu}\equiv x^{\mu}+i\theta\sigma^{\mu}\bar{\theta}$. The lowest component of $\textbf{X}_{nl}$ is eliminated by the nilpotent constraints $\textbf{X}_{nl}^{2}=0$ and is no longer an independent complex scalar field. This leaves only the Goldstino field $\psi_{X}$ and an independent auxiliary field $F_{X}$ in $\textbf{X}_{nl}$ superfield. The Goldstino field $\psi_{X}$, which is the fermionic component of $\textbf{X}_{nl}$, eventually mixes with the gravitino to give it a mass through the super-Higgs mechanism. The lowest component of $\textbf{X}_{nl}$, which is denoted as ${X}_{nl}$ for the Goldstino bilinear $\psi_{X}\psi_{X}$}, obviously satisfies ${X}_{nl}^{2}=0$ due to its anti-commuting Grassman variable nature. The chiral superfield $\textbf{X}_{nl}$ can be seen as the long-distance expression of the microscopic superfield X that satisfies $\bar{D}^{\dot{\alpha}}{\cal J}_{\alpha\dot{\alpha}}=D_{\alpha}\textbf{X}$ for Ferrara-Zumino multiplet ${\cal J}_{\alpha\dot{\alpha}}$ KS , which can be used even when SUSY is spontaneously broken and can be followed consistently along the renormalization group flow down to the deep IR. As noted in Aparicio:2015psl , for the anti-D3 brane in the KKLT Kachru:2003aw scenario, the representation in terms of $\textbf{X}_{nl}$ is very convenient and enough to capture all the anti-D3 brane degrees of freedoms since it allows to treat the effect of anti-D3 brane in terms of standard supergravity couplings of matter and moduli superfields to the nilpotent Goldstino superfield. From the perspective of low energy effective field theory, regardless of the UV origin of SUSY breaking, such a non-linear realization scheme is also convenient for discussions on modulus stabilization and vacuum uplifting.
@@
${\cal L}\supseteq\int d^{2}\Theta 2{\cal{E}}\left[\frac{3}{8}\left(\overline{\cal{D}}^{2}-8{\cal R}\right)\,e^{-K/3}+W\right]+\hbox{h.c.}\,$, (29)
@@
[…]
Integrating out the auxillary field $F_{X}$ gives $F_{X}=-(W_{\textbf{X}_{nl}})^{*}=-f_{0}$. To accommodate the modular flavor symmetry, the Kahler potential and superpotential needs to be adapted to
@@
$W$ $\supseteq$ $\left(f_{0}\textbf{X}_{nl}\left[\eta(\tau)\right]^{2k_{X}}+c_{1}^{3}\right)\frac{H(\tau)}{[\eta(\tau)]^{6}}.$ (31)
To ensure the modular invariance of the superpotential, the modular weight $k_{X}$ should satisfy $k_{X}\in 12\mathbb{Z}$. The supergravity scalar potential can be derived from
@@
$K^{{\alpha}\bar{\beta}}$ $=$ $\left(\begin{array}[]{cc}\frac{(2\Im\tau)^{2}}{3}\left[1-\frac{k_{X}}{3}\frac{{X}^{*}_{nl}{X}_{nl}}{(2\Im\tau)^{k_{X}}}\right]&\frac{i(2\Im\tau)k_{X}{X}_{nl}}{3}\\ -\frac{i(2\Im\tau)k_{X}{X}^{*}_{nl}}{3}&(2\Im\tau)^{k_{X}}\left[1+\frac{k_{X}^{2}({X}^{*}_{nl}{X}_{nl})}{3(2\Im\tau)^{k_{X}}}\right]\end{array}\right)~.$ (36)
Here the indices $\alpha,\beta\sim(\tau,{X}_{nl})$ and the nilpotency of ${X}_{nl}$ is used in the calculation.
From the general expression of the SUGRA scalar potential in (32), the scalar potential for the modulus field $\tau$ and $X_{nl}$ can be calculated as
@@
- $\left.3\left|\tilde{c}_{M}\right|^{2}\left|{H(\tau)}\right|^{2}\right\}~,$
@@
$\tilde{W}_{T}$ $\equiv$ $k_{X}f_{0}X_{nl}\left[\eta(\tau)\right]^{2k_{X}}\left[\frac{i}{2\pi}\hat{G}_{2}(\tau,\tau^{*})+\frac{i}{2\Im\tau}+\frac{ic_{1}^{3}X_{nl}^{*}}{f_{0}(2\Im\tau)^{k_{X}+1}\left[\eta(\tau)\right]^{2k_{X}}}\right]~,$
$\tilde{c}_{M}$ $\equiv$ $f_{0}X_{nl}\left[\eta(\tau)\right]^{2k_{X}}+c_{1}^{3}~,$
$\tilde{W}_{X}$ $\equiv$ $f_{0}\left[\eta(\tau)\right]^{2k_{X}}\left(1+\frac{|X_{nl}|^{2}}{(2\Im\tau)^{k_{X}}}\right)+\frac{{X}^{*}_{nl}c_{1}^{3}}{(2\Im\tau)^{k_{X}}}.$ (38)
The constraint $X_{nl}^{2}=0$ for the lowest component of $\textbf{X}_{nl}$ is used to eliminate some of the high order $X_{nl}$ terms. We also use the following expression for the derivative of Dedekind $\eta$ function
$\frac{\eta^{\prime}(\tau)}{\eta(\tau)}=\frac{i}{4\pi}G_{2}(\tau)~,$ (39)